Solve the differential equation of the following exact differential equation. Show complete solution.


(3x2y - 6x)dx + (x3 + 2y)dy = 0


1
Expert's answer
2021-10-14T18:28:27-0400

"\displaystyle
\text{Let $M(x,y) = 3x^2y-6x$ }\\
\int (3x^2y-6x)dx = x^3y-3x^2+h(y)\\
\text{Differentiating and comparing to $N(x,y) = x^3+2y$, we have that}\\
h'(y) =2y \implies h(y) = y^2\\
\text{Hence the solution to the given exact equation is}\\
x^3y-3x^2+y^2=C"